# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v14

@@ 5.5 Social Changes
Table 6 presents the results for the sex ratio and average household size. Columns 1–2 show the results for the 1930 sample, whereas Columns 3–4 show the results for the 1935 sample. The estimated coefficients in Columns 1 and 2 indicate that municipalities with coal mines experienced an increase in the sex ratio and a decline in the average household size. The estimated magnitudes are economically meaningful. Both estimates ($0.12$ and $0.54$) account for more than one standard deviation in the sex ratio and the average household size, respectively. This confirms that the influx of miners, many of whom were single males, led to a disproportionate sex ratio and smaller household sizes. [^46] This tendency in the mining area is in line with the phenomenon observed in the urban areas in prewar Japan. The urban areas had a relatively higher sex ratio than rural localities due to the internal migration of male workers (Ito 1990, p.243–247). Columns 3–4, however, show much smaller estimates for the 1935 sample. The estimates for the sex ratio and average household size are $0.07$ and $0.49$. While both estimates are still statistically significant, the large declines in magnitude are relevant to the reverting fertility after the regulation (Section 5.3). The increased fertility reverses the overall sex ratio because the sex ratio at birth is generally balanced at that time. It also increases the average household size of existing households, attenuating the impacts of single immigrants.
@@ 5.6 Robustness
The robustness of the results is ensured in the following ways. First, the cluster-robust variance-covariance matrix estimator was used in Online Appendix D.3. This estimator provides a systematically larger estimate for the standard error by allowing for within-cluster dependencies. Despite this, the results are materially similar, meaning that the baseline estimates are robust to the potential influences of the local-scale spatial correlations. Second, the regional heterogeneity in the endowment of male miners is taken into account. The coal mines in Hokkaidō (northernmost island) were less likely to use female miners in the pits, given the nature of the labor supply of male miners. The results are materially similar if we include an indicator variable for Hokkaidō prefecture in the regressions (Online Appendix D.7). Third, an alternative identification strategy is employed. As explained in the data section, the nature of the assignments makes it difficult to employ the estimation strategy utilizing within variations of each unit. Nevertheless, the case study is provided using the DID approach for the subsample of the Tōhoku region (northeastern region). The results are reasonable to the main findings (Online Appendix D.8). Finally, the validity of the results on coal mines is assessed by comparing the results for heavy metal mines. Since coal mining requires many more workers than heavy metal mining because coal is a bulky mineral, the estimated magnitudes for the heavy metal mines shall be much smaller. It is confirmed that the results are reasonable to such expectation (Online Appendix D.9).
@@ 6 Conclusion
The innovation of labor-saving technologies occurred with the renewal of traditional extraction methodology in the coal mining sector during interwar Japan. These technology shocks induced the revision of labor regulations, leading to institutional change concerning the female mines. In this study, the impact of coal mines on the regional economy during historical industrialization in interwar Japan is examined to investigate how the institutional change influenced the effect.
@@
It is found that the coal mines increased the local population, and the estimated magnitude is greater in the sample after the full enforcement of the revised labor regulations. Evidence from the analyses using the register-based vital statistics indicates that the regulations had indeed forced female miners to exit the labor market and form families. Before the full enforcement of the regulation, the occupational hazards for female miners substantially increased the early-life mortality via the biological mortality selection mechanism in utero. The institutional change, however, further led to improvements in early-life mortality.
@@ A.2 Female Miners in Coal Mine
By the end of the 19th century, many mines in the Chikuhō coalfields had finished mechanizing the coal haulage processes. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, this mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312).
Figure A.1a shows a photograph of miners at the mine face (kiriha) in the 1910s. A male skilled miner (saki-yama) extracted coal, whereas a female miner (ato-yama) brought the coal to the surface using a bamboo basket called sura. Figure A.1b shows that ato-yama tied a rope around her body to prevent the heavy sura from slipping off in the steep pits. The environment in the pit was generally poor. Moreover, accidents caused by falling rocks occurred frequently. [^47] Available statistics indicate that the average number of accidents per year that occurred in the pits between 1917 and 1926, causing injuries and deaths of miners, was 142,595, of which 61,112 were from falling rocks. (Japan Mining Association, 1928, p. table.6). Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred 311 times per year on average between 1917 and 1926 (Japan Mining Association 1928, p. table.6)..
@@
The introduction of new mining technologies, such as coal cutters, was gradually adopted in the 1920s. The introduction of these technologies was accompanied by a transition from the traditional room-and-pillar mining method to the longwall method, which is more suitable for mechanical coal extraction. [^48] Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of larger machinery (Nishinarita 1985, p. 91). The number of miners, both male and female, declined during this transition. However, this did not entirely change the labor patterns of female miners. Approximately one in five miners was a woman in 1930 (Table 1), and they still worked inside the pits as ato-yama. [^49] See also Tanaka (1984, chapter. 4) for a comprehensive summary of the relationships between mechanization process and changes in labor patterns in the pits. In short, introducing coal conveyors was the primary driving force for replacing the in-pit works of the female mines. This provided a strong incentive for mining companies to exclude female miners from the pits (Tanaka and Ogino 1977, p. 79). Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work. The number of female miners in the pits (ato-yama) dropped from 1930 to 1933, when the Revised Miner’s Labor Assistance Regulations were fully enacted in September (Table 1). Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. More than 15,000 women still worked in the mines in the late 1930s, representing approximately 10% of all the miners measured.
@@ A.3 Mechanization and Female Miners
Note: This figure illustrates the raw relationship between the share of female miners and log-transformed horsepower in 82 coal mines in Fukuoka prefecture in 1930. The share of female miners is the number of female miners per 100 total miners. The mean and standard deviation of the share of female miners are $20.1$ and $15.6$, respectively. The horsepower is the total horsepower of the machines using any of steam, electrical, and water power. The mean and standard deviation of the log-transformed horsepower are $7.1$ and $2.0$, respectively. The red line represents the fitted values for a linear regression model. Although the Fukuoka Prefectural Statistics originally contained 90 coal mines, 8 mines were dropped because there were missing values in the horsepower statistics.
A.4 Shares of Coal and GSC Mines
@@
Notes: This figure shows the number of coal and oil (Figure A.3a) and metal mines (Figure A.3b) documented in the AHCs of 1931 and 1936. GSC indicates gold, silver, and copper mines.
Source: Created by the author. Data on mines are from the Association for Harmonious Cooperation (1932, 1937).
Figure A.3a shows the number of coal and oil mines documented in the AHCs in 1931 and 1936. Similarly, Figure A.3b shows the numbers of GSC and other metal mines. As shown, approximately 90 (85)of mines extracting fuels (metals) were coal (GSC) mines. Figure A.3 shows the number of mines. Since there were some cases where a few mines coexist in a municipality, the number of municipalities with mines reported in the main text is smaller than the number of mines.
@@
Since prewar Japan was a paternalistic society under the Old Civil Code, the most plausible setting for housework productivity at that time would be $\kappa_{m}=0$. Under this setting, inequality 17 indicates that $\frac{\partial b}{\partial w_{f}}$ is negative under the following condition:
$\footnotesize{\begin{split}\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}.\end{split}}$ (18)
This suggests that female wages negatively correlate with fertility when adequate childcare services are unavailable in local markets.
@@ C.1 Mine Deposit
Data on the location and type of mines are obtained from the official reports published by the Association for Harmonious Cooperation (1932, 1937). They document all mines with more than $50$ miners in October 1931 and October 1936. The documents list three types of mines: metal, coal, and oil. [^50] The metal mine includes minerals such as gold, silver, copper, blister copper, pig iron, steel, iron sulfide, lead, zinc, bismuth, arsenious acid, tin, mercury, chromium, manganese, and sulfide minerals. A smelting mine is also included in the metal mine category. Coal mines include a few mines mining lignite. Oil mines include oil, crude oil, gas, volatile oil, kerosene, light oil, machine oil, and heavy oil. In these documents, the mine’s location is indicated by the name of the municipality. Hence, the name of the municipality was used while geocoding to match the mine dataset with the datasets of the other variables. The shapefile used for geocoding was obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism (available at: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html (accessed on 23rd February 2021). Several municipalities were consolidated between the years 1930 and 1938. Thus, the municipality boundaries defined in 1938 were used and aggregate statistics for consolidated municipalities. A minimal number of municipalities were divided; these were excluded from the original dataset. To check all consolidations and divisions of municipalities, Zenkoku shichōsonmei hensen sōran (Municipal Autonomy Research Association 2006), was used because it is the most comprehensive book on the history of municipalities’ names.
@@
The initial dataset included both coal and heavy metal (gold, silver, and copper) mines. [^51] This is because the most representative heavy metals at that time in the Japanese mining sector were gold, silver, and copper (Online Appendix D.9). Therefore, to prepare the analytical samples, the duplications were excluded to better identify the impact of coal mines. Panel A of Table C.1 summarizes duplications in the coal mine sample. Columns (1) and (2) show the figures based on the 1931 (1936) AHC, which is used to match with the 1930 (1935) Population Census and 1933 (1938) Vital Statistics datasets. Initially, there are $11,151$ municipalities in columns (1) and (2); note that the municipality boundaries are defined based on the 1938 boundaries. The number of municipalities within 30km of a coal mine is $1,147$ and $1,386$ in columns (1) and (2), respectively. The number of duplications in the treatment groups based on the $5$ km radius definition is 00 and $2$ in columns (1) and (2), respectively. Similarly, the number of duplications in the control groups based on the $5$ km radius definition is $7$ and $20$ in columns (1) and (2), respectively. Following the trimming step shown in Figure C.1a, the analytical coal mine samples are $1,140$ ($=1,147-7$) and 1,364 ($=1,386-22$), respectively. Panel B of Table C.1 summarizes the duplications in the GSC mine sample. Following the same procedure as in Figure C.1b, the GSC mine samples used in the robustness analysis are $2,498$ ($=2,510-22$) and $3,293$ ($=3,337-44$), respectively.
@@
The red (blue) circles in Figure C.2 indicate the coal (GSC) mine locations from the AHC. Figure 2 illustrates the spatial distribution of the coal mine sample by year and group. Figure 3 shows the distribution in Chikuhō coalfield in Kyūshū in detail.
@@
Source: Created by the author. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).

### v15

@@ 5.5 Social Changes
Table 6 presents the results for the sex ratio and average household size. Columns 1–2 show the results for the 1930 sample, whereas Columns 3–4 show the results for the 1935 sample. The estimated coefficients in Columns 1 and 2 indicate that municipalities with coal mines experienced an increase in the sex ratio and a decline in the average household size. The estimated magnitudes are economically meaningful. Both estimates ($0.12$ and $0.54$) account for more than one standard deviation in the sex ratio and the average household size, respectively. This confirms that the influx of miners, many of whom were single males, led to a disproportionate sex ratio and smaller household sizes. [^50] This tendency in the mining area is in line with the phenomenon observed in the urban areas in prewar Japan. The urban areas had a relatively higher sex ratio than rural localities due to the internal migration of male workers (Ito 1990, p.243–247). Columns 3–4, however, show similar but slightly smaller estimates for the 1935 sample. [^51] See also Online Appendix D.4 for further evidence from the exercise using the panel structure of the dataset. The estimates for the sex ratio and average household size are $0.07$ and $0.49$. While both estimates are still statistically significant, the declines in magnitude are relevant to the reverting fertility after the regulation (Section 5.3). The increased fertility reverses the overall sex ratio because the sex ratio at birth is generally balanced at that time. It also increases the average household size of existing households, attenuating the impacts of single immigrants.
@@ 5.6 Robustness
The robustness of the results is ensured in the following ways. First, the cluster-robust variance-covariance matrix estimator was used in Online Appendix D.3. This estimator provides a systematically larger estimate for the standard error by allowing for within-cluster dependencies. Despite this, the results are materially similar, meaning that the baseline estimates are robust to the potential influences of the local-scale spatial correlations. Second, the regional heterogeneity in the endowment of male miners is taken into account. The coal mines in Hokkaidō (northernmost island) were less likely to use female miners in the pits, given the nature of the labor supply of male miners. The results are materially similar if we include an indicator variable for Hokkaidō prefecture in the regressions (Online Appendix D.8). Third, an alternative identification strategy is employed. As explained in the data section, the nature of the assignments makes it difficult to employ the estimation strategy utilizing within variations of each unit. Nevertheless, the case study is provided using the DID approach for the subsample of the Tōhoku region (northeastern region). The results are reasonable to the main findings (Online Appendix D.9). In Online Appendix D.4, I further consider a model to test the effects of the labor regulation. Specifically, I consider a difference-in-differences type specification including the interaction term between the treatment group and post-enforcement period dummies. Although the estimates from this type of specification shall be attenuated (Section 5.1), the results obtained are reasonable for my baseline findings. Finally, the validity of the results on coal mines is assessed by comparing the results for heavy metal mines. Since coal mining requires many more workers than heavy metal mining because coal is a bulky mineral, the estimated magnitudes for the heavy metal mines shall be much smaller. It is confirmed that the results are reasonable to such expectation (Online Appendix D.10).
@@ 6 Conclusion
The innovation of labor-saving technologies occurred with the renewal of traditional extraction methodology in the coal mining sector during interwar Japan. These technology shocks induced the revision of labor regulations, leading to institutional change concerning the female mines. In this study, the impact of coal mines on the regional economy during historical industrialization in interwar Japan is examined to investigate how the institutional change was associated with the effect.
@@
It is found that the coal mines increased the local population, and the estimated magnitude is greater in the sample after the full enforcement of the revised labor regulations. Evidence from the analyses using the register-based vital statistics indicates that the regulations had indeed forced female miners to exit the labor market and form families. Before the full enforcement of the regulation, the occupational hazards for female miners increased the early-life mortality via the biological mortality selection mechanism in utero. The institutional change, however, further led to improvements in early-life mortality.
@@ A.2 Female Miners in Coal Mine
By the end of the 19th century, many mines in the Chikuhō coalfields had finished mechanizing the coal haulage processes. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, this mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312). Figure A.1a shows a photograph of miners at the mine face (kiriha) in the 1910s. A male skilled miner (saki-yama) extracted coal, whereas a female miner (ato-yama) brought the coal to the surface using a bamboo basket (sura). Figure A.1b shows that a female miner tied a rope around her body to prevent the heavy sura from slipping off in the steep pits. The environment in the pit was generally poor. Moreover, accidents caused by falling rocks occurred frequently. [^52] Available statistics indicate that the average number of accidents per year that occurred in the pits between 1917 and 1926, causing injuries and deaths of miners, was 142,595, of which 61,112 were from falling rocks. (Japan Mining Association, 1928, p. table.6). Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred 311 times per year on average between 1917 and 1926 (Japan Mining Association 1928, p. table.6)..
(a) Langwall Method
(b) Mechanical Coal Extraction
Figure A.2: Longwall Method, Coal Cutter, and Conveyor
Notes: Figure A.2a shows the system at the coal face used in the single-sided longwall method. This system includes a coal cutter (row III), positioned in front of the coal face (row IV), and a conveyor located behind the coal face (row II). The figure also shows a wagon in the main shaft on the right side. Figure A.2b shows the labor pattern in the coal face under the longwall method, where a male minder uses a scop to transfer coal extracted by the coal cutter to the conveyor behind him.
Sources: Mikawa (1939, p. 274) and Chikuhō Coal Mining Association (1935, p. 88).
The first International Labor Conference in 1919 motivated the government to improve harsh working conditions, especially for female miners. In the early 1920s, the government consulted with mining companies to revise the Miners’ Labor Assistance Regulation of 1916 to prevent females from working in coal mines. Initially, the companies opposed this proposal. However, a shift in mining techniques altered their perspective. By the late 1920s, many coal mines had transitioned from the traditional room-and-pillar method to the longwall method. The longwall method, compared with the room-and-pillar method, offers a more extensive space for coal extraction, which accommodates larger machinery such as coal cutters and conveyors, as shown in Figure A.2a. This provided a strong incentive for mining companies to exclude female miners from the pits (Tanaka and Ogino 1977, p. 79; Nishinarita 1985, p. 91). [^53] See also Tanaka (1984, Chapter 4) for a comprehensive summary of the relationships between the mechanization process and the changes in the labor patterns in the pits. Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited certain types of underground and late-night work, with a grace period for enforcement that lasted until September 1933. During this period of mechanization from the late 1920s to the early 1930s, the number of mines employing both male and female workers decreased. However, the apparent reduction was observed in the female miners who worked in the pits (Table 1). The photograph in Figure A.2b clearly shows how the work of coal extraction at the coal face changed with the introduction of mechanical coal extraction using the longwall method. Additionally, the photographs in Figure A.3 show that the conveyors transported the extracted coal more efficiently than the miners could
(a) Wooden Baskets
(b) Conveyor
Figure A.3: Changes in the Method of Transporting Extracted Coal
from the Coal Face to the Main Shaft
Notes: Figure A.3a shows the traditional coal mining method, where miners use a wooden basket (sura) and must navigate through small, narrow tunnels to reach the main shaft. The same figure also shows the mechanical coal mining method, where a conveyor automatically deposits the extracted coal into a wagon.
Sources: Chikuhō Coal Mining Association (1935, p. 35).
Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. More than 15,000 women still worked in the mines in the late 1930s, representing approximately 10% of all the miners measured.
@@ A.3 Mechanization and Female Miners
Notes: This figure illustrates the raw relationship between the share of female miners and log-transformed horsepower in 82 coal mines in Fukuoka prefecture in 1930. The share of female miners is the number of female miners per 100 total miners. The mean and standard deviation of the share of female miners are $20.1$ and $15.6$, respectively. The horsepower is the total horsepower of the machines using any of steam, electrical, and water power. The mean and standard deviation of the log-transformed horsepower are $7.1$ and $2.0$, respectively. The red line represents the fitted values for a linear regression model. Although the Fukuoka Prefectural Statistics originally contained 90 coal mines, 8 mines were dropped because there were missing values in the horsepower statistics.
@@
A.4 Age Distribution of Female Miners
Figure A.5: Age Distribution of Female Workers in the Coal Mining Sector, as Measured in the 1920 and 1930 Population Censuses
Notes: This figure shows the age distribution of female workers in the mining sector, as recorded in the 1920 Population Census (red) and the 1930 Population Census (green). The percentage of female workers represents the number of female workers per 100 total workers. In the 1920 Population Census, individuals who reported their principal occupation (hongyō) are classified as workers. In the 1930 Population Census, individuals who reported their industry are classified as workers. I test the equality of the age distributions using the Pearson chi-squared statistic (Pearson 1900). The hypothesis of equal distribution is not rejected, with a $p$-value of $0.832$. Data on occupations from the 1925 and 1935 Population Censuses are not available.
Sources: Created by the author. Data are obtained from the Statistics Bureau of the Cabinet (1929, pp. 206–207; 1935b, pp. 160–161).
Figure A.5 compares the age distributions of female workers in the coal mining sector as recorded in the 1920 and 1930 Population Censuses. The 1930 distribution shows a slight shift to the right, which may reflect the effect of the revised labor regulation established in 1928 that prohibited most in-pit work by female and child miners. However, the overall age distribution remains similar to that of the 1920 Population Census. The Pearson chi-squared test does not reject the null hypothesis of equal distribution, with $p$-value of $0.832$. Although occupation statistics are unavailable from the 1935 Population Census, these results suggest that the prohibition of most in-pit work by females and children likely did not affect the age distribution among female miners. Moreover, the proportion of married women reported in the 1920 and 1930 censuses is similar, at over 70% (Statistics Bureau of the Cabinet 1929, pp. 206–207; 1935b, pp. 160–161), suggesting that the productive capacity among females in the mining area might have been stable around 1930.
A.5 Shares of Coal and GSC Mines
@@
Notes: This figure shows the number of coal and oil (Figure A.6a) and metal mines (Figure A.6b) documented in the CSs of 1931 and 1936. GSC indicates gold, silver, and copper mines.
Source: Created by the author. Data on mines are from the Cooperative Society (1932, 1937).
Figure A.6a shows the number of coal and oil mines documented in the CSs in 1931 and 1936. Similarly, Figure A.6b shows the numbers of GSC and other metal mines. As shown, approximately 90 (85)of mines extracting fuels (metals) were coal (GSC) mines. Figure A.6 shows the number of mines. Since there were some cases where a few mines coexist in a municipality, the number of municipalities with mines reported in the main text is smaller than the number of mines.
@@
The inequality 17 indicates that $\frac{\partial b}{\partial w_{f}}$ is negative under the following condition:
$\footnotesize{\begin{split}\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}-w_{m}^{\frac{\rho-1}{\rho}}\kappa_{m}^{\frac{1}{\rho}}.\end{split}}$ (18)
Since prewar Japan was a paternalistic society under the Old Civil Code, the most plausible setting for housework productivity at that time would be $\kappa_{m}\simeq 0$. This means that the second term in 18 is practically negligible. [^54] Roughly, under a specific case of $\kappa_{m}=0$, $\frac{\partial b}{\partial w_{f}}$ is negative under $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$. This suggests that female wages negatively correlate with fertility when adequate childcare services are unavailable in local markets.
@@ C.1 Mine Deposit
Data on the location and type of mines are obtained from the official reports published by the Cooperative Society (1932, 1937). They document all mines with more than $50$ miners in October 1931 and October 1936. The documents list three types of mines: metal, coal, and oil. [^55] The metal mine includes minerals such as gold, silver, copper, blister copper, pig iron, steel, iron sulfide, lead, zinc, bismuth, arsenious acid, tin, mercury, chromium, manganese, and sulfide minerals. A smelting mine is also included in the metal mine category. Coal mines include a few mines mining lignite. Oil mines include oil, crude oil, gas, volatile oil, kerosene, light oil, machine oil, and heavy oil. In these documents, the mine’s location is indicated by the name of the municipality. Hence, the name of the municipality was used while geocoding to match the mine dataset with the datasets of the other variables. The shapefile used for geocoding was obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism (available at: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html (accessed on 23rd February 2021). Several municipalities were consolidated between the years 1930 and 1938. Thus, the municipality boundaries defined in 1938 were used and aggregate statistics for consolidated municipalities. A minimal number of municipalities were divided; these were excluded from the original dataset. To check all consolidations and divisions of municipalities, Zenkoku shichōsonmei hensen sōran (Municipal Autonomy Research Association 2006), was used because it is the most comprehensive book on the history of municipalities’ names.
@@
The initial dataset included both coal and heavy metal (gold, silver, and copper) mines. [^56] This is because the most representative heavy metals at that time in the Japanese mining sector were gold, silver, and copper (Online Appendix D.10). Therefore, to prepare the analytical samples, the duplications were excluded to better identify the impact of coal mines. Panel A of Table C.1 summarizes duplications in the coal mine sample. Columns (1) and (2) show the figures based on the 1931 (1936) CS, which is used to match with the 1930 (1935) Population Census and 1933 (1938) Vital Statistics datasets. Initially, there are $11,151$ municipalities in columns (1) and (2); note that the municipality boundaries are defined based on the 1938 boundaries. The number of municipalities within 30km of a coal mine is $1,147$ and $1,386$ in columns (1) and (2), respectively. The number of duplications in the treatment groups based on the $5$ km radius definition is 00 and $2$ in columns (1) and (2), respectively. Similarly, the number of duplications in the control groups based on the $5$ km radius definition is $7$ and $20$ in columns (1) and (2), respectively. Following the trimming step shown in Figure C.1a, the analytical coal mine samples are $1,140$ ($=1,147-7$) and 1,364 ($=1,386-22$), respectively. Panel B of Table C.1 summarizes the duplications in the GSC mine sample. Following the same procedure as in Figure C.1b, the GSC mine samples used in the robustness analysis are $2,498$ ($=2,510-22$) and $3,293$ ($=3,337-44$), respectively.
@@
The red (blue) circles in Figure C.2 indicate the coal (GSC) mine locations from the CS. Figure 2 illustrates the spatial distribution of the coal mine sample by year and group. Figure 3 shows the distribution in Chikuhō coalfield in Kyūshū in detail.
@@
Source: Created by the author. Data on the location of mines are from the Cooperative Society (1932, 1937).